OpenStudy (delirium):

I will medal and fan. Just help me out, for Pete's sake...whoever in the h*ll Pete is... Write 10 + 6 + 2 + (-2) + (-6) + (-10) + (-14) in sigma notation.

OpenStudy (delirium):

I can convert the Sigma notation to a series. Not the other way around, though.

OpenStudy (mrnood):

What is the common difference?

OpenStudy (delirium):

4

OpenStudy (delirium):

(-)4

OpenStudy (mrnood):

OK and the starting value is 10 so 10 10 + 1(-4) 10 + 2(-4) . . 10 + n(-4)

OpenStudy (delirium):

Right, OK

OpenStudy (mrnood):

note = n starts at ZERO in my sequence.... normally you would start from 1 so last term is (n-1)(-4)

OpenStudy (delirium):

OK, gotch'ya, had to read it a couple of times to process

OpenStudy (delirium):

So...hold on...

OpenStudy (delirium):

\[\Sigma \sum_{\infty}^{n = 1}(10 + n(-4))\]

OpenStudy (delirium):

yea, nay?

OpenStudy (delirium):

Oh, wait, the infinity symbol should be on top, and 'n=1' should be on the bottom...

OpenStudy (delirium):

But you get the point....

OpenStudy (delirium):

\[\sum_{n=1}^{\infty}(10 + n(-4))\]

OpenStudy (mrnood):

no - it does not say it's an infinite series and the first term is 10 so you need to look at my last post...

OpenStudy (mrnood):

(if you put n=1 in your expression what do you get for the first term?)

OpenStudy (delirium):

-40

OpenStudy (delirium):

So, negative 40 goes on top?

OpenStudy (delirium):

(Aren't all series infinite?)

OpenStudy (mrnood):

whoooooa you are NEARLY correct first - it is not an infinite series so you have a definite limit on top - not infinity second if you put n=1 in you get 6 so what is the correct expression?

OpenStudy (delirium):

Oh..duh, why did I multiply ten by -4... OK, then 6 would go on top

OpenStudy (mrnood):

no 1+2 = sum from n=1 to n=2 of n

OpenStudy (delirium):

Oh

OpenStudy (delirium):

Wait... would I just add all the number of terms in the series?

OpenStudy (mrnood):

but you are confusing the terms of the series with the enumeration of n how many terms in the series to get to -14? then look at the valu of the first term, and correct your expression

OpenStudy (mrnood):